# Streetscapes: Large-scale Consistent Street View Generation Using Autoregressive Video Diffusion

### v1

@@ Abstract.
Figure 1. Generating Streetscapes. We design a system that generates Streetscapes, i.e., consistent street views along long-range paths through large-scale synthesised urban scenes. Given a) a street map and its corresponding height map, our system can generate: b) high-quality street views spanning a long camera path with layout controlled by the map, with the ability to generate c) different geographic styles as well as d) different weather conditions or times of day, both controlled by text prompts. We recommend the reader to visit our project page at boyangdeng.com/streetscapes to view our results in videos.
@@ 2. Related Work
Diffusion models have emerged as state-of-the-art generative models for 2D images (Ho et al., 2020a; Song et al., 2020b; Rombach et al., 2022a). These models have also been extended to the 3D domain. For example, some strategies (Nichol et al., 2022; Jun and Nichol, 2023; Shue et al., 2023; Gupta et al., 2023a; Ntavelis et al., 2023) use direct 3D supervision. The quality and diversity of their results, however, is far from that achieved for 2D image generation. DreamFusion (Poole et al., 2022) and related works enable text-to-3D generation by combining a neural scene representation with the image priors of pre-trained 2D diffusion models  (Lin et al., 2023a; Chen et al., 2023a; Wang et al., 2022; Shi et al., 2023; Gupta et al., 2023a). These methods successfully generate 3D objects of moderately high quality. Image-conditioned 3D generation has also become a popular approach (Watson et al., 2022; Liu et al., 2023; Chan et al., 2023; Gu et al., 2023; Hong et al., 2023), with the latest methods in this category (Hong et al., 2023) resolving flicker and other view-inconsistent artifacts that plagued earlier approaches. Based on the idea of training 3D diffusion models directly on image data (Karnewar et al., 2023; Chen et al., 2023b; Shen et al., 2023b), state-of-the-art results for 3D generation have recently been achieved by techniques that solve the 3D generation problem using 2D diffusion models with 3D-aware denoisers (Anciukevičius et al., 2023; Xu et al., 2024).
@@
Video diffusion models  (Singer et al., 2022; Blattmann et al., 2023b; Blattmann et al., 2023a; Li et al., 2023; Guo et al., 2023b; Villegas et al., 2022) have recently emerged as data-driven approaches to training foundation models that can generate short video clips of impressive quality and consistency. ControlNets can be used to condition the video generation process on depth maps or human poses (Hu et al., 2023; Guo et al., 2023a; Xu et al., 2023).
@@ 3.2. Layout-conditioned Scene Generation
We build our framework on top of a pretrained Latent Diffusion Model (Rombach et al., 2022b). Given a noise feature map $z_{T}$ along with a text prompt, this model generates a feature map $z_{0}$ by iteratively denoising $z_{T}$ using Eq. 2. The generated feature map is then decoded into a high-quality image $x_{0}$ using decoder $\mathcal{D},x_{0}=\mathcal{D}(z_{0})$. This decoder, along with a corresponding encoder $\mathcal{E},x\approx\mathcal{D}\left(\mathcal{E}(x)\right)$, enables the diffusion model to operate in a low-resolution feature space, rather than in image space, for computational efficiency. Although not specialized for street-view generation, this foundation model is trained on generic text-to-image generation tasks. We do not further modify its parameters, and so we naturally inherit text control from this foundation model, even for scenarios that are scarce or absent in our data, e.g., streets full of snow.
@@ 3.3. Autoregressive Video Diffusion
Inspired by imputation techniques that were used in time-series analysis (Tashiro et al., 2021), image inpainting (Song et al., 2020c), and monocular depth estimation (Saxena et al., 2023), we propose a temporal imputation method to enable consistent long-range autoregressive video generation for generating Streetscapes.
@@ 4.1. Experimental Details
We base our system on a pretrained text-to-image latent diffusion model (Rombach et al., 2022b). To enable two-frame generation, we insert motion modules as proposed in AnimateDiff (Guo et al., 2023b), but change their $16$-frame formulation to a $2$-frame one. We further add a ControlNet (Zhang et al., 2023) to enable layout and camera pose control. In contrast to the standard single-image ControlNet, we form a two-frame ControlNet by also cloning the motion modules. Hence, our ControlNet contains both spatial and temporal modules. We use the diffusion noise prediction objective from DDPM (Ho et al., 2020b) to train our model with learning rate of $1e^{-5}$ and batch size $256$ for $1.5$M iterations. Note we keep the base diffusion model U-Net intact in training, so as to preserve the text control from the base model. Our model is trained using image data from Google Street View. [^1] Google Maps Street View images used with permission from Google. For two-frame training, we group pairs of $512\times 512$ images captured within $80$s and $5$ meters.
@@ Appendix D Challenges with Geographic Data
Figure 9. Misalignment of Aerial Data and Ground Images. We show $2$ representative examples of misalignment of aerial data (height maps) and the ground images in our dataset. By overlaying the disparity maps rendered using the height maps over the ground image, we notice unneglectable misalignment. Such misalignment includes the offsets near occlusion boundary that we hypothesise to be a systematic error due to the precision of aerial captures; and misalignment of scene content, due to the discrepancy of ever-changing streets shown in ground images for major cities such as New York City and the static height maps available that correspond to only one moment in time. This fact requires our system to be robust to data noise.
@@ Appendix E Ablation Study on Autoregressive Video Diffusion
Figure 11. Impact of Resample and WarpInit. In this experiment, we validate the impact of Resample and WarpInit techniques (see Sec. 3.3) in temporal imputation. Given the source view on the left, we use different variants of temporal imputation, one without the Resample (Ours - Resample), one without the WarpInit (Ours - WarpInit), and the full temporal imputation, to autoregressively generate a target view, for a few samples using different random seeds. Compared to the reference real target view image, we find that our full temporal imputation produces most similar samples, indicating the best consistency in near-range view generation.
@@ Appendix F More Training Details
We base our system on a pretrained text-to-image latent diffusion model (Rombach et al., 2022b). To enable two-frame generation, we first insert motion modules as proposed in AnimateDiff (Guo et al., 2023b), but change their $16$-frame formulation to a $2$- or $4$-frame one. We conduct a two-stage training procedure for our diffusion model. First, we train only the motion modules but keep parameters of the base diffusion model U-Net frozen. We also use per-location text in our training. For each image, we associate it with the text “[location] street” where “[location]” is the city name for the image, e.g., “London”. After training the motion modules, we train a ControlNet (Zhang et al., 2023) for layout and camera pose control. while training the ControlNet, we keep both the base diffusion model and the base motion modules frozen. For both training stages, we use the standard diffusion noise prediction objective from DDPM (Ho et al., 2020b). We train the motion modules with a learning rate of $1e^{-5}$ and batch size at $256$ for $1$M iterations. We train the ControlNet with a learning rate of $1e^{-5}$ and batch size of $256$ for $500$k iterations.

### v2

@@ Abstract.
Figure 1. Generating Streetscapes. We design a system that generates Streetscapes, i.e., consistent street views along long-range paths through large-scale synthesised urban scenes. Given a) a street map and its corresponding height map, our system can generate: b) high-quality street views spanning a long camera path with layout controlled by the map, with the ability to generate c) different geographic styles as well as d) different weather conditions or times of day, both controlled by text prompts. We recommend the reader to visit our project page at boyangdeng.com/streetscapes to view our results as videos.
@@ 2. Related Work
Diffusion models have emerged as state-of-the-art generative models for 2D images (Ho et al., 2020; Song et al., 2020b; Rombach et al., 2022). These models have also been extended to the 3D domain. For example, some strategies (Nichol et al., 2022; Jun and Nichol, 2023; Shue et al., 2023; Gupta et al., 2023a; Ntavelis et al., 2023) use direct 3D supervision. The quality and diversity of their results, however, is far from that achieved for 2D image generation. DreamFusion (Poole et al., 2023) and related works enable text-to-3D generation by combining a neural scene representation with the image priors of pre-trained 2D diffusion models  (Lin et al., 2023a; Chen et al., 2023a; Wang et al., 2022; Shi et al., 2023; Gupta et al., 2023a). These methods successfully generate 3D objects of moderately high quality. Image-conditioned 3D generation has also become a popular approach (Watson et al., 2022; Liu et al., 2023; Chan et al., 2023; Gu et al., 2023; Hong et al., 2023), with the latest methods in this category (Hong et al., 2023) resolving flicker and other view-inconsistent artifacts that plagued earlier approaches. Based on the idea of training 3D diffusion models directly on image data (Karnewar et al., 2023; Chen et al., 2023b; Shen et al., 2023b), state-of-the-art results for 3D generation have recently been achieved by techniques that solve the 3D generation problem using 2D diffusion models with 3D-aware denoisers (Anciukevičius et al., 2023; Xu et al., 2024).
@@
Video diffusion models  (Singer et al., 2022; Ho et al., 2022; Blattmann et al., 2023b; Blattmann et al., 2023a; Li et al., 2023; Guo et al., 2023b; Villegas et al., 2022) have recently emerged as data-driven approaches to training foundation models that can generate short video clips of impressive quality and consistency. ControlNets can be used to condition the video generation process on depth maps or human poses (Hu et al., 2023; Guo et al., 2023a; Xu et al., 2023).
@@ 3.2. Layout-conditioned Scene Generation
We build our framework on top of a pretrained Latent Diffusion Model (Rombach et al., 2022). Given a noise feature map $z_{T}$ along with a text prompt, this model generates a feature map $z_{0}$ by iteratively denoising $z_{T}$ using Eq. 2. The generated feature map is then decoded into a high-quality image $x_{0}$ using decoder $\mathcal{D},x_{0}=\mathcal{D}(z_{0})$. This decoder, along with a corresponding encoder $\mathcal{E},x\approx\mathcal{D}\left(\mathcal{E}(x)\right)$, enables the diffusion model to operate in a low-resolution feature space, rather than in image space, for computational efficiency. Although not specialized for street-view generation, this foundation model is trained on generic text-to-image generation tasks. We do not further modify its parameters, and so we naturally inherit text control from this foundation model, even for scenarios that are scarce or absent in our data, e.g., streets full of snow.
@@ 3.3. Autoregressive Video Diffusion
Inspired by imputation techniques that were used in time-series analysis (Tashiro et al., 2021), image inpainting (Song et al., 2020b), monocular depth estimation (Saxena et al., 2023), and text-conditioned generic video generation (Ho et al., 2022), we propose a temporal imputation method to enable consistent long-range autoregressive video generation for generating Streetscapes.
@@ 4.1. Experimental Details
We base our system on a pretrained text-to-image latent diffusion model (Rombach et al., 2022). To enable two-frame generation, we insert motion modules as proposed in AnimateDiff (Guo et al., 2023b), but change their $16$-frame formulation to a $2$-frame one. We further add a ControlNet (Zhang et al., 2023) to enable layout and camera pose control. In contrast to the standard single-image ControlNet, we form a two-frame ControlNet by also cloning the motion modules. Hence, our ControlNet contains both spatial and temporal modules. We use the diffusion noise prediction objective from DDPM (Ho et al., 2020) to train our model with learning rate of $1e^{-5}$ and batch size $256$ for $1.5$M iterations. Note we keep the base diffusion model U-Net intact in training, so as to preserve the text control from the base model. Our model is trained using image data from Google Street View. [^1] Google Maps Street View images used with permission from Google. For two-frame training, we group pairs of $512\times 512$ images captured within $80$s and $5$ meters.
@@ Appendix D Challenges with Geographic Data
Figure 9. Misalignment of Aerial Data and Ground Images. We show two representative examples of misalignment of aerial data (height maps) and the ground images in our dataset. By overlaying disparity maps rendered from the height maps atop the corresponding ground-level RGB views, we observe noticeable misalignment. Such misalignment includes the offsets near occlusion boundaries that may be attributable to systematic error due to the precision of aerial captures. We also see misalignment of scene content, due to the discrepancy between the the static height maps we use and the ever-changing street scenes characteristic of vibrant cities like NYC. These discrepancies and misalignments necessitate that our system be robust to data noise.
@@ Appendix E Ablation Study on Autoregressive Video Diffusion
Figure 11. Impact of Resample and WarpInit. In this experiment, we validate the impact of Resample and WarpInit techniques (see Sec. 3.3) in temporal imputation. Given the source view on the left, we use different variants of temporal imputation, one without Resample (Ours - Resample), one without WarpInit (Ours - WarpInit), and the full temporal imputation, to autoregressively generate a target view, for a few samples using different random seeds. Compared to the reference real target view image, we find that our full temporal imputation produces most similar samples, indicating the best consistency in near-range view generation.
@@ Appendix F More Training Details
We base our system on a pretrained text-to-image latent diffusion model (Rombach et al., 2022). To enable two-frame generation, we first insert motion modules as proposed in AnimateDiff (Guo et al., 2023b), but change their $16$-frame formulation to a $2$- or $4$-frame one. We conduct a two-stage training procedure for our diffusion model. First, we train only the motion modules but keep parameters of the base diffusion model U-Net frozen. We also use per-location text in our training. For each image, we associate it with the text “[location] street” where “[location]” is the city name for the image, e.g., “London”. After training the motion modules, we train a ControlNet (Zhang et al., 2023) for layout and camera pose control. while training the ControlNet, we keep both the base diffusion model and the base motion modules frozen. For both training stages, we use the standard diffusion noise prediction objective from DDPM (Ho et al., 2020). We train the motion modules with a learning rate of $1e^{-5}$ and batch size at $256$ for $1$M iterations. We train the ControlNet with a learning rate of $1e^{-5}$ and batch size of $256$ for $500$k iterations.